40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 30 02:13:23 2017, gautam, Update, General, DRMI locking attempt BS_OLmods.pdf
    Reply  Fri Jun 30 11:50:17 2017, gautam, Update, General, DRMI locking attempt ITMX_glitchy.png
       Reply  Fri Jun 30 22:28:27 2017, gautam, Update, General, DRMI re-locked DRMI_relocked.pngMICH_OL.pdf
          Reply  Sat Jul 1 14:27:00 2017, Koji, Update, General, DRMI re-locked 
Message ID: 13090     Entry time: Fri Jun 30 11:50:17 2017     In reply to: 13088     Reply to this: 13093
Author: gautam 
Type: Update 
Category: General 
Subject: DRMI locking attempt 

Seems like the problem is actually with ITMX - the attached DV plots are for ITMX with just local damping loops on (no OLs), LR seems to be suspect.

I'm going to go squish cables and the usual sat. box voodoo, hopefully that settles it.

Quote:

Summary:

I attempted to re-lock the DRMI and try and realize some of the noise improvements we have identified. Summary elog, details to follow.

  1. Locked arms, ran ASS, centered OLs on ITMs and BS on their respective QPDs.
  2. Looked into changing the BS Oplev loop shape to match that of the ITMs - it looks like the analog electronics that take the QPD signals in for the BS Oplev is a little different, the 800Hz poles are absent. But I thought I had managed to do this successfully in that the error signal suppression improved and it didn't look like the performance of the modified loop was worse anywhere except possibly at the stack resonance of ~3Hz --- see Attachment #1 (will be rotated later). The TRX spectra before and after this modification also didn't raise any red flags.
  3. Re-aligned PRM - went to the AS table and centered beam on all REFL PDs
  4. Locked PRMI on carrier, ran MICH and AS dither alignment. PRC angular feedforward also seemed to work well.
  5. Re-aligned SRM, looked for DRMI locks - there was a brief lock of a couple of seconds, but after this, the BS behaviour changed dramatically.

Basically after this point, I was unable to repeat stuff I did earlier in the evening just a couple of hours ago. The single arm locks catch quickly, and seem stable over the hour timescale, but when I run the X arm dither, the BS PITCH loop starts to oscillate at ~0.1 Hz. Moreover, I am unable to acquire PRMI carrier lock. I must have changed a setting somewhere that I am not catching right now (although I've scripted most of these things for repeatability, so I am at a loss what I'm missing indecision). The only change I can think of is that I changed the BS Oplev loop shape. But I went back into the filter file archives and restored these to their original configuration. Hopefully I'll have better luck figuring this out tomorrow.

 

Attachment 1: ITMX_glitchy.png  36 kB  Uploaded Fri Jun 30 12:51:36 2017  | Show | Show all
ELOG V3.1.3-